A banker invested a portion of $10,000 at 5% interest per year and the remainder at 4% interest per year. If the banker received $470 in interest after 1 year, how much was invested at each rate?
Answer by mananth(16949)

Part I 5.00% per annum
Part II 4.00% per annum

x+y=10000 ------------------------1
5.00%x+ 4.00%y =$470.00
Multiply by 100
5x+4y=$47,000.00 --------2

Multiply (1) by -5
we get

-5x-5y =-50000.00

Add this to (2)

-y=-$3,000.00
y=$3,000.00 investment at 4.00%
Balance $7,000.00 investment at 5.00%